OpenStudy (anonymous):

Let y'(x) + y(x)g'(x) = g(x)g'(x), y(0) = 0, x Î R, where ƒ'(x) denotes dƒ(x)/dx and g(x) is a given non-constant differentiable function on R with g(0) = g(2) = 0. Then the value of y(2) is

OpenStudy (anonymous):

we need to multiply both sides by g(x)

OpenStudy (anonymous):

(gy)'=g^2g'

OpenStudy (anonymous):

hmm i don't know let me think some more

OpenStudy (anonymous):

hmm i don't know let me think some more

OpenStudy (anonymous):

i thing i got divide both sides by g' so we have y'/g'+y=g so we want (1/g')'=1 => 1/g'=x =>g'=1/x so then we have (xy)'=g => xy=g+C y=g/x+C/x

OpenStudy (anonymous):

if g'=1/x then g=lnx+C

OpenStudy (anonymous):

but this isn't defined at 0 so again i don't know lol

OpenStudy (anonymous):

i made a mistake above we have (xy)'=g right! =>(xy)'=1/x xy=ln(x)+C => y=ln(x)/x+C/x but still this doesn't help

OpenStudy (anonymous):

try determining c

OpenStudy (anonymous):

oh we are given y(0)=0 but still we can't plug in ln(0) doesn't exit 1/0 also doesn't exist

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

i think i got it so we have y'+yg'=gg' if we multiply both sides by v we get vy'+vg'y=vgg' so we want this v'=vg' so we have dv/dx=v*dg/dx dv=v*dg 1/v dv=dg integrating both sides gives ln|v|=g+k let k=0 so v=e^g so we have e^g*y'+e^g*g'y=e^g*g' which we can write as (e^gy)'=(e^g)' integrate both sides e^gy=e^g+C y=1+C/e^g ....

OpenStudy (anonymous):

i have no clue :(

OpenStudy (anonymous):

i have no clue :(